I don't know what to do first in solving this word problem. It's really confusing me.
"You jog at a speed of 6 miles per hour for 20 minutes, and you walk at a speed of 4 miles per hour for t minutes. If your combined jogging and walking distance is 3 miles, for how long did you walk?"

Answer by stanbon(48568)

"You jog at a speed of 6 miles per hour for 20 minutes, and you walk at a speed of 4 miles per hour for t minutes. If your combined jogging and walking distance is 3 miles, for how long did you walk?
----------------
Jogging distance:
rate = 6 mph ; time= (1/3) hr. ; distance =rt= 6(1/3)= 2 mi.
---------
Walking distance:
rate = 4 mph ; time = t/60 hr. ; distance = rt = 4(t/60)=(t/15)
----------
EQUATION:
jog dist + walk dist = 3 miles
2 + (t/15) = 3
t/15 = 1
t= 15 minutes
You walked for 15 minutes.
Cheers,
Stan H.